# Climate in a Bottle: Towards a Generative Foundation Model for the Kilometer-Scale Global Atmosphere

### v1

@@ Abstract
AI emulators offer a path to compressing, boosting limited ensembles, and improving the latency of interacting with petabyte-scale climate prediction data. However, prevailing auto-regressive paradigms offer limited flexibility, and are challenging to train on climate time horizons due to drifts, instabilities and component-coupling challenges. Conditionally generative models offer an appealing alternative. In this context we demonstrate a generative diffusion-based framework—Climate in a Bottle (cBottle)—for emulating global km-scale climate simulations and reanalysis on the equal-area HEALPix grid. cBottle consists of two model stages: a globally-trained coarse-resolution image generator that generates 100km (50k-pixel) fields given monthly average sea surface temperatures and solar conditioning, followed by a locally-trained 16x super-resolution stage that generates 5km (12.5M-pixel) fields; global super-resolution is made affordable using an overlapping patch-based multi-diffusion. Overall, cBottle shows promise as an emulator across a battery of climate model diagnostics, including diurnal-to-seasonal scale variability, large-scale modes of variability, tropical cyclone statistics, and trends of climate change and weather extremes. Moreover, cBottle is a step towards a foundation model, by bridging multiple data modalities (reanalysis and simulation) with corresponding utility beyond emulation to tasks such as zero-shot bias correction, climate downscaling, and channel in-filling. The code is available at https://github.com/NVlabs/cBottle.
@@ 1 Introduction
Knowing more, and planning better for the consequences of climate change are driving tremendous advances in the realms of climate simulation and Earth observation. With exascale computing and energy-efficient computational systems, petabytes of kilometer-scale climate data are being generated. Digital-twin efforts like the Destination Earth initiative (Bauer et al., 2021) and the Earth Virtualization Engines (EVE) project (Stevens et al., 2024) envision a world where high-fidelity global climate projections with local granularity are available to all with a degree of access and interactivity that provides a scientifically grounded basis for action. Further, digital twins aim to efficiently expose the data to new methods of analysis and proactive information production that will enable stakeholders to construct and interact with their own climate scenarios.
@@
Easy access and low-latency interactivity, however, are nearly impossible with the current paradigm of petabyte-scale datasets generated and hosted by a handful of climate centers. Even if the access problem is unthrottled by moving data around, interacting with petabytes of data is virtually impossible. Generative artificial intelligence (AI) is poised to disrupt this paradigm by fully leveraging state-of-the-art architectures and algorithms optimized for performance on accelerated computing hardware.
For our purposes, we define a generative foundation model as a self-supervised climate simulator capable of producing realistic-looking samples that are climate-faithful measured by a set of diagnostics, and can be used to examine what-if scenarios not explicitly built into the model. Of course, climate science has long had such foundation models in the form of physics-based general circulation models. However, advances in AI are already fundamentally changing how weather and climate are simulated due to their ease of use, enhanced computational performance, and improved accuracy. Specifically, emulators of weather have proven to be skillful and orders-of-magnitude faster than traditional numerical simulation-based approaches to simulating the weather across spatial and temporal scales (Pathak et al., 2022; Lam et al., 2023; Bi et al., 2023; Price et al., 2024; Bodnar et al., 2024). Similarly, AI is proving to be efficient and faithful at emulating climate models and climate data, evaluated across a range of metrics on a wide variety of tasks (Watt-Meyer et al., 2024; Chapman et al., 2025; Cresswell-Clay et al., 2024; Watson-Parris et al., 2022). A majority of these developments have resorted to the auto-regressive modeling paradigm, much like numerical weather and climate models that step forward in time evolving the state of the system modeled by governing systems of equations. Masked training and representation learning have also been proposed (Lessig et al., 2023).
@@
Diffusion modeling (Ho et al., 2020; Song et al., 2020a; Karras et al., 2022) offers an alternative to the auto-regressive time-stepping approach by learning—and sampling—from a distribution of states. This is especially attractive for rapid analyses and informatics that do not require the full temporal evolution of the system. By the same token, they are not susceptible to instabilities or drifts (i.e., hallucinations) inherent to auto-regressive models, though admittedly many auto-regressive emulators perform surprisingly well for large numbers of steps (Watt-Meyer et al., 2024; Cresswell-Clay et al., 2024). Further, diffusion models can be controlled by the user in a variety of ways, using guidance (Song et al., 2020a) to generate conditional distributions. Generative diffusion modeling has had tremendous impact across disciplines ranging from computer vision to science, and enjoys widespread adoption across the industry. Notably, it has been used in weather and climate for ensemble generation and uncertainty quantification (Li et al., 2024), probabilistic forecasting (Price et al., 2024), super-resolution (or downscaling) (Mardani et al., 2025), and data-assimilation (Rozet and Louppe, 2023; Manshausen et al., 2024).
@@
A less explored yet promising direction is its application towards learning a compressed representation of massive climate datasets that generates, on demand, climate states steered by user-specified conditions and scenarios. Early examples included training deterministic emulators to predict the mean-state response of climate models to various forcings (Watson-Parris et al., 2022). More recent efforts have extended this approach to stochastic emulators in the diffusion framework (Bassetti et al., 2024; Quilcaille et al., 2022) as well as traditional dimensionality reduction approaches (Wang et al., 2025). But these approaches have been limited to a small number of channels (typically surface temperature) at a coarse spatial resolution.
@@
We develop a generative diffusion-based framework that we call “Climate in a bottle” (cBottle) to emulate multimodal climate data. cBottle uses a cascaded approach (Ho et al., 2021) to generating km-scale samples: a coarse-resolution generator conditioned on user-specified inputs (time of day, time of year, and monthly-averaged sea surface temperature (SST)) and a 16x super-resolution module. Both steps use generative diffusion models, and either can be used independently, enabling generation of coarse-resolution samples and/or super-resolution to km-scale.
cBottle has a few key properties:
@@
    Multi-modality and domain transfer: Bridges ERA5 and ICON km-scale datasets, enables generation of km-scale climate states that exploits the information content contained across ERA5 and ICON
@@
    Versatility across a few useful tasks: Bias-correction, channel in-filling, downscaling
@@
    Climate faithfulness: Generates realistic climatology, climate variability across timescales, large-scale modes, and extreme weather statistics
@@
    Data efficiency: Can be trained on as little as 4 weeks of simulation output
@@
    Extreme compression: Encapsulates km-scale climate model and reanalysis outputs into a few GB of neural network weights, and offers 256x compression ratio per channel
@@
    Speed and scale: Low-latency generation of km-scale samples and scalability to higher resolutions and more modalities
These properties of cBottle offer potential pathways towards: (i) solving the data and informatics challenges of km-scale climate data; and (ii) building a generative foundation model for an interactive digital twin.
Mean sea-level pressure (generated) (ground truth)
@@
Surface air temperature (generated) (ground truth)
@@
Infrared(generated) (ground truth)
Refer to caption Refer to caption
Visible light (generated) (ground truth)
Refer to caption Refer to caption
Surface precipitation (generated) (ground truth)
Refer to caption Refer to caption
@@
Figure 1: High resolution (HPX1024) samples from the ICON data (right) and synthesized by cBottle given sea surface temperature and time conditionings (left). Note these are not identically paired. For each field the same colormap is used, but to highlight the visual similarity we do not include color-bars. More quantiative comparisons will follow.

### v2

@@ Abstract
Climate modeling is reaching unprecedented resolution, producing petabytes of climate data [1]. AI climate model emulators offer a path to computationally cheap analysis of this data, enabling new scientific insight and scenario planning [2]. Recent advances have shown promise in faithfully emulating climate data [3, 4, 5, 6]. However, prevailing auto-regressive paradigms are challenging to train on climate time horizons due to drifts, instabilities and component-coupling challenges, they are hard to scale to high resolution, and pose the challenge of having to sift through troves of output to identify rare climate extremes of outsize interest. Here, we present Climate in a Bottle (cBottle), a generative diffusion-based framework [7, 8, 9] emulating global 5 km climate simulations and reanalysis on the HEALPix grid. cBottle directly samples from the full distribution of atmospheric states without requiring a previous time step and therefore auto-regressive rollout, as well as the first to reach this global resolution (12.5M-pixel). cBottle consists of two model stages [10]: a global coarse-resolution generator conditioned on sea surface temperatures and solar position, followed by a patch-based 16x super-resolution stage. cBottle passes a battery of climate model tests, including diurnal-to-seasonal scale variability, large-scale modes of variability, tropical cyclone statistics, and trends of climate change and weather extremes. Moreover, cBottle is a step towards a foundation model: It bridges multiple data modalities (reanalysis and simulation), and enables zero-shot bias correction, climate downscaling, and data in-filling. It also enables a new form of human interactivity by leveraging guided diffusion [8]. For example, users can request samples of extreme weather located over regions of interest – we demonstrate this by training a tropical cyclone (TC) classifier alongside the generator, guiding towards TC states, and obtaining physically credible samples. We anticipate this to kickstart developments of guidance methods for a wide array of user queries and new ways of interacting with climate data.
@@ 1 Introduction
Policymakers, infrastructure planners, the insurance industry and the public rely on accurate and high-resolution information about the present and possible future climate. This need is driving tremendous advances in climate simulation and Earth observation technology. With exascale computing, energy-efficient computational systems now generate petabytes of kilometer-scale climate data. Digital-twin efforts like the Destination Earth initiative [1] and the Earth Virtualization Engines (EVE) project [2] envision such high-fidelity physical climate projections as underpinning next-generation climate informatics infrastructure, which should eventually enable a degree of access and interactivity to allow widespread climate-informed decision making.
@@
Easy access and low-latency interactivity, however, are nearly impossible with the current paradigm of petabyte-scale datasets generated and hosted by a handful of climate centers. Even if the access problem is unthrottled by moving data, interacting with petabytes of data is computationally expensive.
Recent advances in AI to replace the simulation task do not solve this challenge. While AI has proved efficient and faithful at emulating models and climate data evaluated across a range of metrics on a wide variety of tasks [3, 4, 5, 6], most of these developments have resorted to the auto-regressive modeling paradigm, much like numerical weather and climate models that step forward in time evolving the state of the system modeled by governing systems of equations. Such systems have not been scaled to the km-scale global resolution envisioned by EVE. Even if this were possible it would not solve the informatics challenge of nimbly guiding users to samples or statistics of interest, such as rare extreme weather events impacting a particular region. That is, autoregressive AI emulators share a common challenge with physical climate models – the need to sift through troves of simulation output to uncover information of interest.
@@
We propose that generative artificial intelligence (AI) for sampling offers a solution to both the data volume and interactivity problems by fully leveraging state-of-the-art algorithms optimized for performance on accelerated computing hardware. For our purposes, we define a generative foundation model as a self-supervised climate simulator capable of producing realistic-looking samples that are climate-faithful measured by a set of diagnostics, and can be used to examine what-if scenarios not explicitly built into the model. Specifically, diffusion modeling [7, 8, 9] offers an alternative to the auto-regressive time-stepping approach by learning—and sampling—from a distribution of states. This is especially attractive for rapid analyses that do not require the full temporal evolution of the system. By the same token, they are not susceptible to instabilities or drifts inherent to auto-regressive models, though admittedly many auto-regressive emulators perform surprisingly well for large numbers of steps [3, 5]. Further, diffusion models can be controlled by the user in a variety of ways, using guidance [8] to generate conditional distributions.
Generative diffusion modeling has been used in weather and climate for ensemble generation and uncertainty quantification [11], probabilistic forecasting [12], super-resolution (or downscaling) [13], and data-assimilation [14, 15].
@@
A less explored yet promising direction is its application towards learning a compressed representation of massive climate datasets that generates, on demand, climate states steered by user-specified conditions and scenarios. Early examples included training deterministic emulators to predict the mean-state response of climate models to various forcings [6]. More recent efforts have extended this approach to stochastic emulators in the diffusion framework [16, 17] as well as traditional dimensionality reduction approaches [18]. But these approaches have been limited to a small number of channels (typically surface temperature). And, as with current AI simulators, none have been scaled to the ambitious km-scale resolution of next-generation climate predictions such as proposed by EVE.
In this context, we present “Climate in a bottle” (cBottle), a first-of-its-kind generative diffusion-based framework, faithfully emulating multimodal high-resolution climate data. cBottle offers pathways towards: (i) solving the data and informatics challenges of km-scale climate data; and (ii) building a generative foundation model for an interactive Earth digital twin.
2 cBottle
@@
cBottle generates km-scale samples from the climate distribution using a cascaded approach [10] as shown in Fig. 1b): A coarse-resolution generator conditioned on user-specified inputs (time of day, time of year, and monthly-averaged sea surface temperature (SST)) and a 16x super-resolution module capable of generating over ten million pixels globally (see Fig. 2 for a visualization of the high-resolution output) for dozens of channels (see Table 2). Both steps use generative diffusion models, and either can be used independently, enabling generation of coarse-resolution samples and/or super-resolution to km-scale. Special care was required to train the model to produce the largest modes of variability without over-fitting to them. See Methods for details of the model formulation. Briefly, cBottle’s key properties are:
@@
    Multi-modality and domain transfer: Trained on reanalysis data and 5-km global atmospheric simulations with a convection-resolving climate model, enables generation of km-scale climate states that exploits the information content contained in both modalities. Specifically, we use the ECMWF Reanalysis v5 (ERA5) [19] —a 40-year observationally-constrained dataset available at coarse resolution, and a 5-year simulation with the ICON global 5-km model [20].
@@
    Versatility across several useful tasks: removing bias in climate simulations, diagnosing some fields (e.g. precipitation) given others (e.g. winds), or down-scaling coarse-resolution climate data to km-scale.
@@
    Climate faithfulness: Generates realistic climatology, climate variability across timescales, large-scale modes, and extreme weather statistics,
@@
    Data efficiency: The super-resolution model can be trained on as little as 4 weeks of simulation output, reducing the need for very long simulations with expensive convection-resolving models.
@@
    Extreme distillation: Encapsulates km-scale climate model and reanalysis outputs into a few GB of neural network weights, and offers 256x compression ratio per channel.
@@
    Speed and scale: Low-latency generation of km-scale samples and scalability to higher resolutions and more modalities.
  7. 7.
    Interactivity: Can be guided flexibly to yield extreme events like Tropical Cyclones in a specific location on demand.
Figure 1: Climate in a Bottle: A generative foundation model for kilometer-scale climate simulation. (a) Training phase: Petabytes of high-resolution climate data from ICON global cloud resolving simulations and ERA5 reanalysis are compressed into a compact neural network representation ($\sim$ few GB). (b) Inference phase: The trained model generates realistic 5km-resolution climate states from minimal conditional inputs (time, location, sea surface temperature) through a two-stage process: Coarse global generation (100km) followed by patch-based super-resolution. (c) Applications: The foundation model enables diverse climate science applications including interactive what-if scenarios, bias correction between datasets, channel in-filling for missing variables, statistical downscaling, multimodal learning across data sources, and high-fidelity climate emulation—all achievable in minutes rather than hours and accessible without supercomputing resources.
Surface air temperature (tas) Ground truth Generated
@@
Infrared (rlut) Ground truth Generated
@@
Figure 2: High resolution (HPX1024) samples from the ICON data (left) and synthesized by cBottle given sea surface temperature and time conditionings (right). Note these are not identically paired but use the same input SST condition for 2025-03-16T15:00:00. For each field the same colormap is used, but to highlight the visual similarity we do not include color-bars. More quantitative comparisons will follow. See Figure SI1 for more fields.
3 Results
Below, the acronyms HPX64 (HPX1024) refer to the low-resolution 100-km (high-resolution 5-km) components of cBottle (see Section 5.2).
3.1 ICON in a bottle
Generation of ICON states at HPX1024 resolution – nearly 13M pixels, a formidable task – is achieved through a two-stage process in which the macroscale generator of the ICON modality at HPX64 is fed through a patch-based super-resolution module (see Methods), which we validate below. Figure 2 shows a comparison of (unpaired) samples of ICON and cBottle end-to-end-generated surface temperature and infrared radiation—these are visually similar, and the generated samples show what fine-grained detail cBottle can achieve. While we have chosen to highlight the end-to-end synthesis task – where high-dimensional ICON weather states are generated from low-dimensional climate-controlling inputs – it is not the only use-case. The compression application, where coarsened ICON output are input to the super-resolution model also is relevant as a means of compressing large global storm-resolving simulations. Validation of super-resolution in this context is provided in Section B.4.1 where we show cBottle successfully super-resolves the different length scales of ICON’s convective organization from shallow cloud systems influencing its shortwave radiation channel to deep tropical convection clusters in its precipitation channel (Fig. SI2), and provide quantitative validation of PDFs and spectra (Fig. SI8).
3.2 ERA5 in a Bottle
Realistic conditional samples
To be useful as a climate sampler, cBottle must demonstrate realistic variability across independent samples from phenomena operating on diurnal, synoptic, seasonal, and interannual time scales that, when averaged, produce a realistic climatology. To validate this we perform an Atmosphere Model Intercomparison Project (AMIP) control experiment with prescribed monthly-averaged SSTs [21, 22]. Samples from AMIP experiments are typically produced by integrating physical climate models (or autoregressive AI simulators) forward in time subject to SST boundary conditions. In cBottle, analogous samples can be obtained by simply generating independent samples every hour from 1940 to 2021 (inclusive)—a large-scale inference, requiring 1024$\sim$ 1024∼ 1024 GPU-hours on the H100 GPU, but embarrassingly parallel since each atmospheric state (timestep) is completely independent and only shares the same conditioning.
The following common diagnostics assess the quality of the generated samples and statistics. Additional analyses are presented in Section B.6.
Figure 3: Fidelity to conditioning. Amplitude of the diurnal cycle of precipitation for (a) cBottle and (b) ERA5. DJF-JJA Arctic sea ice concentration computed over the test period for (c) cBottle and (d) ERA5. (e) cBottle’s secular temperature trend during the AMIP period compared to ERA5. (f) User-provided hurricane conditioning. Minima of pressure shown in blue are linked to TCs generated by cBottle using classifier guidance. The input to the inference is the location of the cyclones, shown as turquoise boxes.
Diurnal Cycle
In Figure 3a,b) we compare the amplitude of the diurnal cycle of precipitation in cBottle with ERA5. Although the amplitude is too large in cBottle, there is a realistic spatial structure of the precipitation diurnal cycle amplitude including spatial details interior to tropical continents – and even the halo that surrounds the Maritime Continent from the emanating diurnally phased convection propagation there [23]. Section B.1 provides additional analysis of the diurnal band including the signal-to-noise ratio which confirms cBottle reproduces even small-amplitude signals like the locally detectable subtropical stratocumulus drizzle diurnal cycles (cf. Fig. SI5(d)) and the appropriate land-sea contrast in the surface temperature diurnal cycle (Fig. SI4).
Seasonal cycle and Mean Climate
Figure 3c) shows that cBottle correctly generates the seasonal evolution of Arctic sea ice concentration compared to ERA5 in d) (See Fig. SI14 for the concentrations by season). During boreal summer (JJA) the sea ice margin retreats and its concentration is reduced locally over the Beaufort Sea, Chukchi Sea and Bering Sea – poleward of Alaska and Siberia. These results are evidence that the time-of-year and/or SST-conditioning in cBottle work as intended. The seasonal cycle is similarly realistic for other fields (see Figure SI15). The annual mean climate of the cBottle generated precipitation samples also matches the ERA5 and ICON data (see Figure SI16).
Northern Annular Mode and SST conditioning
We next assess a planetary scale mode of internal atmospheric variability. In the northern hemispheric midlatitudes, the Northern Annular Mode (NAM) is the leading empirical orthogonal function (EOF), which accounts for one-fifth of overall variance in the monthly anomalies of 1000 hPa geopotential height computed from November to April [24]. Figure 4d shows that the daily variability of cBottle’s NAM index is plausible though slightly weak—with a root-mean-square amplitude in 2018 of 0.84 versus 1.02 in ERA5. Moreover, its low frequency envelope is in phase with inter-annual fluctuations in the ERA5 NAM state, which is more variable during the winter months, suggesting learnt associations between the generated NAM state and the SST and day of year inputs to cBottle. Appropriate conditioning on SST is further confirmed by analyzing the El Nino Southern Oscillation (ENSO; see SI Section B.2 and Fig. SI6). We note that this success in NAM generation depended on using our proposed regularized ensemble of experts approach (see Section A.3.2) to avoid severe over-fitting to the training data that is otherwise encountered when a single checkpoint trained to 9.8M images is naively used instead. Figure 4b demonstrates such overfitting as unrealistic temporal coherence between ERA5 and cBottle in the test set; in contrast, our regularized model has a better fair continuous-ranked-probability score (CRPS) [25] in the validation set and a smaller discrepancy between training and validation scores.
Figure 4: Internal variability. (a,c) Tropical cyclone occurrence probability derived from ERA5 and cBottle (1980–2018), presented as heatmaps normalized per $km^{2}$. (b,d) The Northern Annular Mode (NAM) index for cBottle samples with overfitting (b) and large-noise-level regularization (d). The NAM indices are computed by projecting the daily Z1000 anomalies onto the NAM loading pattern (Fig SI13. For convenience, the CRPS is computed by comparing the 0 UTC from the ground truth index to the 24-member ensemble of cBottle samples taken from ever hour of the same day.
3.3 Interactivity
cBottle being a diffusion model enables the unique capability of oversampling parts of the distribution that are particularly relevant, such as extreme events. We show this using the example of Tropical Cyclones (TCs), which are generated with a realistic climatology in vast AMIP samples (Fig. 4a,b) but nonetheless rare in individual generations. This relates to a principal motivation of cBottle to enable new interactive experiences in climate informatics, such as sidestepping the need for a user to sift through vast datasets to find e.g. extreme events of interest over locations of concern.
To demonstrate user-informed sampling from the distribution of TCs, we train a simple classifier on historical observations from the International Best Track Archive for Climate Stewardship (IBTrACS) [26]. We use the classifier for guidance during inference [27] (see section 5.7 for details), enabling us to steer the output state to be compatible with an input map of desired Tropical Cyclone locations. Fig. 3f shows a sample generated using this technique, with three TCs successfully generated in the locations requested in the conditioning input. This demonstrates that an AI climate state sampler can generate extreme events on-demand.
In our experiments, imposing such conditions in certain locations did not necessarily lead to reduced variability elsewhere, and channels are correlated in a realistic way. Unphysical conditioning, such as guiding towards a TC on the equator, is possible but is somewhat impeded by the model (see SI section B.3).
3.4 Multi-modal inference
We demonstrate how the separate modalities and generative vs. super-resolution modules of cBottle can be composed into novel workflows, providing the beginnings of a foundational versatility. We show three examples tested on a real-world atmospheric-river that occurred during the validation period [28].
Channel infilling
When the channel of interest is not only biased, but damaged, or not available at all in one of the two datasets, we can use cBottle to infill it based on the other dataset. In Fig. 5 e–g) we show infilled radiation channels for the ERA5 sample in a, b), channels which in the original ERA5 data show streaking artifacts and were excluded from the training data. Compared to the ERA5 ground truths in Fig. 5 a, b), cBottle successfully reconstructs plausible ICON fields. This feature can be used to diagnose any given field given arbitrary subsets of the other fields so it removes the need to train combinatorially many diagnostic models to predict certain fields (e.g. precipitation) given others (e.g. free-tropospheric states).
Translating between modalities
The ICON model prefers a liquid-heavy partitioning between cloud water and ice compared to ERA5 (cf Fig. 5a, c). It has nearly double the liquid water as ERA5. cBottle can translate from such an ERA5 sample (a, b) to an ICON sample (c, d) (see Methods). Overall, the location of the macroscale weather phenomena in the translated data is similar, though there are some apparent differences. This ability to translate between datasets with a single foundation model is an exciting potential application, and we leave more detailed analysis to future studies.
Downscaling
To downscale ERA5 data to higher resolution, we first translate the ERA5 sample to the ICON distribution using the coarse-resolution model, and then apply the super-resolution model. The results in Fig.5h shows that the precipitation is concentrated in the same regions as Fig. 5a,b.
In summary, these results demonstrate that cBottle can plausibly reconstruct physical variables and unresolved features throughout the globe. This underscores the potential of cBottle as a valuable tool for data recovery in atmospheric science, particularly in scenarios where observational records are incomplete, corrupted, or degraded by systematic artifacts or high-resolution dynamical-downscaling simulations have not been performed.
Figure 5: Multi-modal inference methods. The original ERA5 data for vertically integrated liquid water (CLLVI) (a) and $dbR=10log(precip)$ (b). (c-d) These fields translated to the ICON distribution. (e-g) The missing radiation channels infilled by cBottle-3d: outgoing shortwave (RSUT), surface downwelling shortwave (RSDS), and outgoing longwave (RLUT) (h) The downscaled dbR for this same example. This is computed by applying cBottle-SR after translation from ERA5 to ICON. These are shown for an atmospheric river example from 2006-07-11 12UTC.
4 Conclusions
We have introduced a generative AI model capable of synthesizing full global atmospheric states at 5 km resolution—over 12.5 million pixels per sample—without requiring autoregressive rollout. This milestone in resolution and sampling fidelity is achieved through a two-stage diffusion framework: a coarse-resolution generator conditioned on minimal, climate-controlling inputs (monthly SST, time of year/day), followed by a patch-based super-resolution module. Together, they enable low-latency, high-fidelity generation of multivariate climate states at the same ambitious resolution of the world’s next-generation operational climate simulation technology.
In addition to resolution, cBottle demonstrates foundational versatility. Trained jointly on reanalysis (ERA5) and global cloud-resolving simulations (ICON), it permits multimodal translation between datasets, channel in-filling for corrupted or missing variables, bias correction and downscaling and zero-shot generalization of its global super-resolution module. This is made possible by a shared latent representation and masked training strategy (see Methods), allowing cBottle to flexibly bridge data modalities and tasks.
Perhaps most importantly, cBottle proposes a new form of interactive climate modeling. By integrating classifier guidance into the diffusion process, users can conditionally generate physically plausible extreme events—such as tropical cyclones—at specified locations. This demonstrates that an AI climate model can sample rare events on-demand, opening the door to targeted scenario generation and new forms of climate risk analysis.
Methodologically, we have introduced diffusion modeling on the HEALPix grid and proposed a noise schedule tuned to the spectral characteristics of climate data. In particular we show (see Methods) that the defaults most often used in natural image generation literature are insufficient for generating extremely large-scale modes like the seasonal and diurnal cycles. Fixing this required two main innovations: i) increasing the range of noise levels sampled during training and inference; and ii) adding appropriate time and position embedding. Specifically, we propose tuning the noise schedule based on spectral analysis of the data instead of expensive hyper-parameter tuning. We have also identified and mitigated over-fitting at high noise levels using an ensemble-of-experts approach; it is logical to speculate that similar issues may explain the lack of calibration in regional down-scaling models [29]. By comparison our super-resolution models trained on large global storm-resolving models simulations are much less prone to over-fitting both owing to the quantity of the training data and the enhanced locality of the super-resolution task compared to global state generation. These perspectives should be helpful to anyone training diffusion models on scientific data.
Additional limitations of our work are important to note. Non-stationary trends are not optimally represented in the current version of cBottle, with its long-term trend only half as big as observed. Similarly, the amplitude of NAM index of the generated samples is 20% too small, suggesting the model under-represents the variance in the largest scales. We attribute both of these weaknesses to the heavy-handed regularization required to prevent the coarse model from over-fitting at large noise levels. Nonetheless the modularity, scalability, and interactivity of cBottle represent a step towards a truly interactive Earth digital twin.